What is the pH of a 1.5x10^-3 M solution of NaOH, a strong base? Show your work.


1
Expert's answer
2022-05-31T13:34:03-0400

1.

First we find pOH . pOH is the negative of the logarithm of the OH ion concentration:

pOH = - log [OH].

Since NaOH => Na+ + OH , the OH concentration is equal to the NaOH concentration. That is 1,5*10–3 mol/l

pOH is the negative of the logarithm of the OH ion concentration.

pOH = - log [OH] = -log (1,5 * 10–3) =

= 3 - log (1,5) = 3 - 0,176 = 2,824


2.

Knowing that the sum of pH and pOH in the solution is 14, we find the pH :

pH = 14 – pOH = 14 – 2,824 = 11,176


ANSWER: pH = 11,176